Opened 10 years ago

Last modified 6 years ago

#415 assigned flight processing

BB11_01, flight day 244a/2011, Brecon Beacons

Reported by: knpa Owned by: adbe
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Processing: general Keywords:
Cc: mggr Other processors:

Description (last modified by knpa)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/uk/BB11_01-2011_244a_Brecon_Beacons

Data arrived from ARSF via network transfer on 02/09/2011

BB11_01 Scientific objective:

BB11_01 Priority: a4h

PI: Shaun Lewis (Brecon Beacons National Park Authority)

Sensors:

Camera (11/10/2011)
Eagle (23/11/2011)
Hawk (23/11/2011)
LiDAR (23/11/2011)

Special note

Atmospheric correction requested. Please check with Ben, Steve or Mike prior to creating a normal delivery.

Change History (33)

comment:1 Changed 10 years ago by mggr

  • Cc mggr added
  • Description modified (diff)

This project is the one that we're going to explore atmospheric correction on. Please check with Ben, Steve or myself before we make a delivery :)

comment:2 Changed 10 years ago by knpa

  • Description modified (diff)

comment:3 Changed 10 years ago by knpa

  • Description modified (diff)

comment:4 Changed 10 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:5 Changed 10 years ago by knpa

  • Owner set to adbe
  • Status changed from new to assigned

adbe has started navigation processing.

comment:6 Changed 10 years ago by adbe

  • Description modified (diff)

I have split the project into two, as the two areas covered are quite far apart:

  • BB11_01-2011_244a_Brecon_Beacons
  • BB11_01-2011_244b_Brecon_Beacons

comment:7 Changed 10 years ago by adbe

Processed navigation

comment:8 Changed 10 years ago by knpa

  • Summary changed from BB11_01, flight day 244/2011, Brecon Beacons to BB11_01, flight day 244a/2011, Brecon Beacons

comment:9 Changed 10 years ago by jaho

Started camera processing.

Removed images 1 - 17 (taken before the flight).

comment:10 Changed 10 years ago by jaho

Camera ready for delivery check.

Starting Lidar processing.

comment:11 Changed 10 years ago by jaho

Lidar ready for delivery check.

comment:12 Changed 10 years ago by knpa

Starting camera delivery check.

comment:13 Changed 10 years ago by knpa

Delivery check complete.

All looks good apart from:

  1. The metadata tiff tags should give the sortie as "a" (because this is 244a).
  2. Filenames also need the flight letter e.g. RCD105-BB11_01-2011244-00075.tif should be RCD105-BB11_01-2011244a-00075.tif. Also, change KML name and names in event file.
  3. First photo in events file has nonsense pitch/roll/heading

comment:14 Changed 10 years ago by adbe

Started hyperspectral processing, need to use a nextmap dem for the mapped level 3.

comment:15 Changed 10 years ago by jaho

Correcting camera for 244a and 244b.

comment:16 Changed 10 years ago by adbe

Created hyperspectral delivery. Note: not included DEM as it is NextMap

comment:17 Changed 10 years ago by jaho

Camera 244a and 244b fixed.

comment:18 Changed 10 years ago by knpa

Starting LiDAR delivery check.

comment:19 Changed 10 years ago by knpa

Delivery check complete.

All looks fine apart from:

  1. Remove logsheet_overview.png from logsheet directory
  2. Line names have a on end in ReadMe but not in logsheet. (e.g. 001a compared to 001)
  3. Delete the .aux.xml files with DEM
  4. Is there a reason SRTM used and not ASTER? Don't worry about it for this delivery but ASTER should normally be used.
  5. Have you checked with PI if they want to wait for vectors? If not, you need to e-mail them. If they want to wait for vectors then you will need to check against these when they arrive and regenerate intensity image so it has the vectors overlaid.

comment:20 Changed 10 years ago by jaho

Waiting for PI's address to send 244a and 244b camera and lidar.

comment:21 Changed 10 years ago by jaho

Camera and LiDAR sent to Shaun Lewis on hard drive 114 (11/10/2011).

comment:22 Changed 10 years ago by benj

ASD Fieldspec and Microtops measurements received from BBNPA and stored in the flight directory under ground_measurements

comment:23 Changed 10 years ago by knpa

Gary says that the PI should be able to get their own vectors via PGA.
Do not wait for vectors to make delivery but deliver with caveat that we have not made this usual check.

comment:24 Changed 10 years ago by jaho

Reprocessed LiDAR flightlines with pitch offset -0.00291419. Created new ASTER dem in place of the old SRTM one. Updated elevation differences. The data seems to look better. New delivery at: delivery/BB11_01-244a-lidar-20111117.

comment:25 Changed 10 years ago by knpa

Doing HS delivery check.

comment:26 Changed 10 years ago by knpa

  1. Need to mention in Readme that we've not checked against vectors.
  2. Might have to unix2dos the bad pixel files?
  3. h6 has some drop frames in middle of line - needs to be mentioned in Readme

Otherwise good, ready to deliver.

comment:27 Changed 10 years ago by adbe

Hyperspectral:

  • Mentioned in ReadMe about vectors and hawk dropped frames.
  • Done unix2dos on .txt files

comment:28 Changed 10 years ago by jaho

Mentioned about vectors in LiDAR ReadMe. Fixed clouds.

comment:29 Changed 10 years ago by adbe

  • Description modified (diff)

Hyperspectral(No atmospheric correction) and LiDAR(correct pitch) delivered to Shaun Lewis on 23/11/2011 on disk 117

comment:30 Changed 10 years ago by knpa

  • Description modified (diff)

Everything has been delivered for this project, however should not be set for archiving as some atmospheric correction work is still being done on it (Peter Land?)

comment:31 Changed 10 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:32 Changed 10 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.04-0.07
2-0.02-0.05
3-0.05-0.13
4-0.03-0.03
5-0.03-0.05
6-0.01-0.02
7-0.06-0.06

comment:33 Changed 6 years ago by dac

Tidied up by moving from processed/delivery to delivery. Hardlinked files into final LiDAR delivery (DEM only) to create complete delivery. Previous LiDAR deliveries moved to delivery/previous_deliveries

Note: See TracTickets for help on using tickets.